# Large language model psychometrics: A systematic review of evaluation, validation, and enhancement

### v2

@@ 8.6 Item response theory
Applying IRT to LLM outputs offers several advantages: more informative benchmarking [Guinet et al., 2024], identification of items that best discriminate among high-performing models [Lalor et al., 2024], and adaptive testing to reduce evaluation costs [Zhuang et al., 2023a, Truong et al., 2025]. Most work to date uses standard IRT models (1PL, 2PL, 3PL), with little exploration of polytomous, hierarchical, or fully multidimensional models, despite the complexity of LLM capabilities. Integrating IRT with generative AI for automated item generation is emerging, but current methods are often too complex for practical use [Jiang et al., 2024a]. Although IRT could standardize measurement of LLMs and humans on a unified scale—even with different test items—this potential is largely untapped.
Systematic use of IRT for bias analysis, such as detecting differential item functioning across model families, is still nascent [He-Yueya et al., 2024]. IRT provides a principled way to uncover biases by analyzing performance across items of varying difficulty and demographic context. For example, if items involving certain groups are disproportionately difficult, this may reveal model or data biases. IRT may also strengthen robustness evaluation by testing model performance on adversarial or edge-case items.
@@ 8.7 From evaluation to enhancement
Evaluation should not only deepen our understanding of models but also drive their improvement. While psychometric research has focused on evaluating, comparing, and interpreting LLMs, applying these insights to model enhancement remains underdeveloped. As noted in § 7, psychometric principles can guide prompt engineering, inference-time control, training data selection, reward design for fine-tuning, and methodological development. Enhancement methods are still nascent. For example, value targets for LLM alignment in Yao et al. [2024], Ye et al. [2025b] are statically defined, limiting the integration of context-dependent value states seen in humans [Skimina et al., 2021]. Leveraging psychometric insights can enable more effective model improvement strategies.
@@ 9 Ethical considerations
Psychometric evaluation and enhancement of LLMs risk amplifying existing biases or enabling manipulation, especially when using unrepresentative data or instruments. Such tools often reflect the values of specific groups, potentially marginalizing others or promoting narrow perspectives [Wang et al., 2025a]. Psychometric insights can also be used to shape LLM behavior for particular interests, raising concerns about fairness and transparency. Researchers should ensure data and instruments are inclusive and representative, and clearly report limitations and potential biases.
@@
LLM Psychometrics often analyzes human-LLM interactions and benchmarks models against human participants, raising privacy and consent concerns. Evaluating LLMs with real user data or human studies requires strong safeguards: informed consent, data anonymization, and transparent data usage. Researchers must also prevent unintended psychological profiling or behavioral analysis of individuals through their LLM interactions.
@@ 10 Conclusion
This paper presents a systematic review of LLM Psychometrics. The integration of psychometric instruments, theories, and principles into LLM evaluation promises to overcome the limitations of traditional AI benchmarks. This approach enables us to more effectively capture the broad, emergent psychological constructs of LLMs, encompassing both personality and cognitive dimensions. Psychometric evaluation methodologies vary in test formats, data and task sources, prompting strategies, model outputs, and scoring methods. They each have distinct strengths, weaknesses, and application scenarios, yet all must adhere to psychometric principles such as reliability, validity, and fairness. Beyond evaluation, psychometrics-inspired techniques enhance LLMs in trait manipulation, safety and alignment, and cognitive capabilities, contributing to the development of more powerful and responsible AI systems.
AI development is increasingly driven by evaluation [Silver and Sutton, 2025, Yao, 2025, AAAI, 2025]. We argue that LLM Psychometrics will play a pivotal role in this shift, introducing new principles, dimensions, techniques, and insights. We hope this review inspires future evaluation paradigms for human-level AI and advances AI psychology for the greater good.

### v3

@@ 8.6 Item response theory
Applying IRT to LLM outputs offers several advantages: more informative benchmarking [guinet2024automated], identification of items that best discriminate among high-performing models [lalor2024item], and adaptive testing to reduce evaluation costs [zhuang2023efficiently, truong2025reliable]. Most work to date uses standard IRT models (1PL, 2PL, 3PL), with little exploration of polytomous, hierarchical, or fully multidimensional models, despite the complexity of LLM capabilities. Integrating IRT with generative AI for automated item generation is emerging, but current methods are often too complex for practical use [jiang2024raising]. Although IRT could standardize measurement of LLMs and humans on a unified scale—even with different test items—this potential is largely untapped.
Systematic use of IRT for bias analysis, such as detecting differential item functioning across model families, is still nascent [he2024psychometric]. IRT provides a principled way to uncover biases by analyzing performance across items of varying difficulty and demographic context. For example, if items involving certain groups are disproportionately difficult, this may reveal model or data biases. IRT may also strengthen robustness evaluation by testing model performance on adversarial or edge-case items.
@@ 8.7 From evaluation to enhancement
Evaluation should not only deepen our understanding of models but also drive their improvement. While psychometric research has focused on evaluating, comparing, and interpreting LLMs, applying these insights to model enhancement remains underdeveloped. As noted in §˜7, psychometric principles can guide prompt engineering, inference-time control, training data selection, reward design for fine-tuning, and methodological development. Enhancement methods are still nascent. For example, value targets for LLM alignment in yao2024value_fulcra, ye2025generative_psycho_lexical are statically defined, limiting the integration of context-dependent value states seen in humans [skimina2021between]. Leveraging psychometric insights can enable more effective model improvement strategies.
8.8 Standardization and norming
While reliability, validity, and fairness have been adapted with varying degrees of success in LLM psychometrics (§˜6), two fundamental psychometric properties remain significantly underdeveloped: standardization and scale (norming). These represent critical gaps that require both theoretical and empirical advancement.
Standardization.
In human psychometric testing, standardization ensures strict, uniform protocols for test administration. However, LLM evaluations lack consensus on standardized inference parameters (temperature, top-p sampling, system prompts) and prompting strategies (§˜5.3). This heterogeneity makes cross-study comparisons problematic and undermines reproducibility. Establishing standardized evaluation protocols (e.g., via a public leaderboard [yao2025value_compass_leaderboard]) is essential for the field to mature toward cumulative, comparable science.
Scale and norming.
Norming in psychometrics establishes reference distributions that render raw scores interpretable. While human test scores derive meaning from comparisons to defined populations (e.g., “90th percentile among college students”), the appropriate reference population for LLMs remains undefined. Current research typically reports raw scores without context, applies human norms (implicitly assuming a construct equivalence that may be invalid, §˜8.4), or compares models within ad-hoc subsets constrained by rapid development and selective sampling. Given that a single model’s profile can shift dramatically based on prompting strategies, and that the population of models is rapidly evolving, the applicability of traditional norming concepts in LLM Psychometrics remains an open question.
@@ 9 Ethical considerations
Psychometric evaluation and enhancement of LLMs risk amplifying existing biases or enabling manipulation, especially when using unrepresentative data or instruments. Such tools often reflect the values of specific groups, potentially marginalizing others or promoting narrow perspectives [wang2025large]. Psychometric insights can also be used to shape LLM behavior for particular interests, raising concerns about fairness and transparency. Researchers should ensure data and instruments are inclusive and representative, and clearly report limitations and potential biases.
@@
LLM Psychometrics routinely benchmarks LLMs against human participants and repurposes large-scale normative datasets—personality inventories, value surveys, moral judgment corpora—collected from participants who consented to research on human psychology, not to their responses serving as reference baselines or training signals for AI systems. This secondary-use consent gap manifests in three ways: (1) normative comparison, where human response distributions define the standard against which LLM outputs are judged; (2) fine-tuning and alignment, where individual-level psychometric responses are used directly as training signal; and (3) instrument repurposing, where items from copyrighted or access-restricted tests are administered to LLMs without licensing consideration. Beyond these data issues, studies involving live human–LLM interactions require informed consent, data anonymization, and transparent usage policies to prevent unintended psychological profiling. We call for field-level norms: reporting the provenance and original consent scope of human comparison data, and preferring open, freely licensed instruments or purpose-built LLM benchmarks where possible.
@@ 10 Conclusion
This paper presents a systematic review of LLM Psychometrics. The integration of psychometric instruments, theories, and principles into LLM evaluation promises to overcome the limitations of traditional AI benchmarks. This approach enables us to more effectively capture the broad, latent psychological constructs of LLMs, encompassing both personality and cognitive dimensions. Psychometric evaluation methodologies vary in test formats, data and task sources, prompting strategies, model outputs, and scoring methods. They each have distinct strengths, weaknesses, and application scenarios, yet all must adhere to psychometric principles such as reliability, validity, and fairness. Beyond evaluation, psychometrics-inspired techniques enhance LLMs in trait manipulation, safety and alignment, and cognitive capabilities, contributing to the development of more powerful and responsible AI systems.
AI development is increasingly driven by evaluation [silver2025welcome, ysymyth_second_half, aaai2025_ai_evaluation]. We argue that LLM Psychometrics will play a pivotal role in this shift, introducing new principles, dimensions, techniques, and insights. We hope this review inspires future evaluation paradigms for human-level AI and advances AI psychology for the greater good.